# MNRAS  template – title goes here

### v2

@@ Abstract
In this work, we present a method for numerically solving the Friedmann equations of modified $f(\mathcal{G})$ gravity in the presence of pressureless matter. This method enables us to predict the redshift behavior of the Hubble expansion rate. To evaluate the credibility of the model, we applied a Bayesian MCMC technique using late-time cosmic observations to impose limitations on the free parameters of the Gauss-Bonnet model. Our results suggest that the $f(\mathcal{G})$ model can reproduce the low-redshift behavior of the standard $\Lambda$CDM model, but there are significant differences at high redshifts, leading to the absence of a standard matter-dominated epoch. We also examined the profiles of cosmographic parameters using the model parameter values from the standard range to verify the intermediate epochs. Our analysis shows that the highly promising $f(\mathcal{G})$ model is a feasible candidate for explaining the current epochs. We presented a dynamical system analysis framework to examine the stability of the model. Our study identified critical points depicting various phases of the Universe and explained the evolutionary epochs. We demonstrated that the model effectively captures the evolution of energy components over cosmic time, supporting its validity as an alternate explanation for the observed acceleration of the Universe.
@@
cosmology: theory, cosmology: dark energy
@@ 1 Introduction
The quest to understand the gravitational force led to the development of General Relativity (GR) by Einstein (1915). This theory revolutionized our understanding of gravity by describing it as the curvature of spacetime caused by the presence of mass and energy (Misner et al., 1973). Although GR has successfully passed numerous experimental tests (Will, 2014) and remains the cornerstone of modern gravitational physics, it struggles to fully explain phenomena such as the accelerated expansion of the Universe (Riess et al., 1998), dark matter (Bertone et al., 2005), and dark energy (DE) (Copeland et al., 2006). In response to these challenges, several modified theories of gravity have been proposed (Clifton et al., 2012).
@@
In this work, we have explored a subclass of the $f(\mathcal{G})$ model to test its viability as an alternative to the standard cosmological paradigm. We have developed a numerical method to predict the redshift behavior of the Hubble expansion rate, and our results suggest that the model can reproduce the low-redshift behavior of the standard $\Lambda$CDM model but has significant differences at high redshifts. The $f(\mathcal{G})$ model is a feasible candidate for explaining the current epochs and effectively captures the evolution of energy components over cosmic time, supporting its validity as an alternative explanation for the observed acceleration of the Universe. We delved into the background cosmological dynamics of the chosen model and evaluated its feasibility using Bayesian analysis supported by Markov Chain Monte Carlo (MCMC) methods applied to late-time cosmic observations, such as Supernovae Ia (Pantheon+) and observational Hubble data (CC sample). We have introduced a dynamical system analysis framework to assess the stability of the model. Our research pinpointed critical points that illustrate different phases of the Universe and elucidated the evolutionary epochs. We have shown that the model effectively represents the changing energy components over cosmic time, supporting its credibility as an alternative explanation for the observed acceleration of the Universe.
This work aims to establish constraints on $f(\mathcal{G})$ cosmology models using CC and Pantheon+ datasets. The paper comprehensively analyses $f(\mathcal{G})$ gravity and uses dynamical system analysis to investigate the stability of the model. The mathematical formalism of $f(\mathcal{G})$ gravity is detailed in Section 2. In Section 3, we use the MCMC method to establish correlations between the $f(\mathcal{G})$ gravity model and observational data to determine the best fits for the model parameters $H_{0}$, $\alpha$, $\beta$, and $m$. Additionally, we present plots of various cosmological parameters such as deceleration, effective equation of state (EoS) $\omega_{\text{eff}}=\frac{p_{\text{eff}}}{\rho_{\text{eff}}}$, Om diagnostic, and the $r-s$ parameter plot, which are essential for understanding the dynamical behavior of the Universe under $f(\mathcal{G})$ gravity. Subsequently, in Section 4, we construct a dynamical system framework to analyze the critical points of the $f(\mathcal{G})$ gravity model. This analysis is crucial for assessing the stability and viability of the model and its alignment with current cosmological observations. Finally, in Section 5, we present the conclusions of our results.
@@ 2 $R+f(\mathcal{G})$ gravity
The parameter for the effective dark energy EoS can be defined as follows:
@@ 2.1 Power-Law $f(\mathcal{G})$ Model
In order to determine the theoretical values of the Hubble rate, we can calculate it by solving equation (8a) numerically. When we assume that matter behaves as a pressureless perfect fluid, we can express the matter density as $\rho_{m}=3H_{0}^{2}\Omega_{\text{m}0}(1+z)^{3}$, with $z$ representing the cosmological redshift (defined as $\frac{a_{0}}{a}=1+z$, where $a_{0}$ represents the scale factor at present and $a$ denotes the scale factor when the light emitted) and $\Omega_{\text{m}0}$ representing the current value of the matter density parameter. Thus, for the particular model we are examining, the first Friedmann equation can be expressed as follows:
@@ 3 Observation with numerical solution
In this part, we will assess the observational feasibility of the model under examination by conducting a Bayesian analysis of the late-time cosmic data. Specifically, we will evaluate the data from the SNe Ia Pantheon+ sample (Brout et al., 2022) and the cosmic chronometers (CC) derived from the observational Hubble data compiled in Ref. (Moresco et al., 2022). We have not assumed that the Hubble and Pantheon datasets are correlated. Rather, we will present our results independently for the CC and Pantheon+ datasets. Utilising these datasets for statistical analysis enables us to obtain reliable results that are not influenced by assumptions of any particular underlying reference model (D’Agostino & Luongo, 2018; D’Agostino, 2019; Lohakare et al., 2023). In the following subsections, we will outline the key characteristics of these measurements and the corresponding likelihood functions.
@@ 3.1 Cosmic Chronometers (CC)
Figure 1: Contour plots show the $1-\sigma$ and $2-\sigma$ uncertainty regions for the variables $H_{0}$, $\Omega_{m0}$, $\alpha$, $\beta$, and $m$. These contours are derived from the CC sample (left panel) and the Pantheon+ data (right panel).
@@ 3.2 Supernovae type Ia (SNe Ia)
The contour plot for the parameters $H_{0}$ and $\Omega_{\text{m}0}$ exhibits a more elliptical shape compared to the other parameter pairs, which display more square or elongated contours. This elliptical shape indicates that there is a relatively weak correlation between $H_{0}$ and $\Omega_{\text{m}0}$, suggesting that the data independently constrains these two parameters. This independence implies that the variations in one parameter do not significantly affect the value of the other, leading to a more symmetrical uncertainty region. In contrast, the square or elongated contours seen in other parameter pairs indicate stronger correlations, where changes in one parameter can be offset by adjustments in another to maintain a similar fit to the data. This strong coupling results in less symmetrical and more stretched uncertainty regions, reflecting the interdependence of these parameters within the $f(\mathcal{G})$ gravity model. In order to evaluate the efficacy of our MCMC analysis, we calculated the associated AIC, $\text{AIC}_{\text{c}}$ and BIC values, which are presented in Table 2. Our findings strongly support the assumed $f(\mathcal{G})$ gravity models when analyzing the datasets. Moreover, we noted that the $f(\mathcal{G})$ model demonstrates greater precision when applied to the Pantheon+ datasets.
@@
Figure 3: Whisker plot depicting the model parameters $H_{0}$, $\Omega_{\text{m}0}$, $\alpha$, $\beta$ and $m$, respectively with their $1-\sigma$ confidence intervals.
@@ 3.3 Cosmological Parameter Evolution
We analyze the evolution of crucial cosmological parameters, including the effective EoS, statefinder, and Om diagnostic parameters, by imposing constraints on model parameters using various observational data. We present a fully general expression for the deceleration parameter, $q=-\ddot{a}/aH^{2}$, as follows:
@@
Figure 4: Graphical representation of the deceleration parameter versus redshift using the constrained coefficients from Fig. 1. The thick line represents the behavior of the deceleration parameter for the $f(\mathcal{G})$ model, while the dashed line shows the deceleration parameter for the $\Lambda$CDM model.
@@
Fig. 4 presents the deceleration parameter as a function of redshift $z$ for the $f(\mathcal{G})$ gravity model, derived from both Hubble data (upper panel) and Pantheon+ data (lower panel). Fig. 4 demonstrate that the restricted values of model parameters derived from the analyzed CC and Pantheon+ data sets indicate a transition of $q$ from positive (indicating early deceleration) in the past to negative (indicating current acceleration) in the present. The present value of deceleration parameter $q_{0}$ is measured to be $-0.527$ and $-0.499$ for the CC and Pantheon+ data, respectively, at the current cosmic epoch, which aligns relatively well with the range of $q_{0}=-0.528^{+0.092}_{-0.088}$ determined by recent observations (Gruber & Luongo, 2014). Recent observations are consistent with this deceleration parameter, and the resulting model indicates a smooth transition from deceleration to acceleration at $z_{t}=0.84$ and $z_{t}=0.82$ for the CC and Pantheon+ datasets, respectively. The derived transition redshift $z_{t}$ aligns with current constraints based on 11 $H(z)$ observations reported by Busca et al. (2013) for redshifts $0.2\leq z\leq 2.3$, $z_{t}=0.74\pm 0.5$ from Farooq & Ratra (2013), $z_{t}=0.7679^{+0.1831}_{-0.1829}$ by Capozziello et al. (2014), and $z_{t}=0.60^{+0.21}_{-0.12}$ by Yang & Gong (2020). The consistency between the curves from Hubble and Pantheon+ data underscores the robustness of the $f(\mathcal{G})$ model in capturing the expansion dynamics of the Universe.
@@
Figure 5: Graphical representation of the EoS parameter versus redshift using the constrained coefficients from Fig. 1. The thick line represents the behavior of the EoS parameter for the $f(\mathcal{G})$ model, while the dashed line shows the EoS parameter for the $\Lambda$CDM model.
@@
The deceleration parameter is one of the key factors that characterize the behavior of the Universe, determining whether it continuously decelerates, accelerates, or undergoes multiple phases of transition. Similarly, energy sources influence the evolution of the Universe through the EoS parameter, defined as $\omega_{\text{DE}}$ is shown in Fig. 5. By calculating the energy density and pressure of DE, as depicted in Fig. 5, we can observe the variations in the effective EoS of DE relative to the redshift variable. The current EoS values for DE, $\omega_{\text{DE}}(z=0)$, are obtained as $-1.018$, $-0.999$ for the CC, Pantheon+, datasets, respectively. These values indicate phantom behavior (at $z<0$) and a trend towards approximately $-1.32$ at late times. The present values of $\omega_{\text{eff}}$ are $-0.684$, $-0.666$ for the CC and Pantheon+ datasets, respectively. Various cosmological studies have also constrained the EoS parameter, including the Supernovae Cosmology Project $\omega_{\text{DE}}=-1.035^{+0.055}_{-0.059}$ (Amanullah et al., 2010), Planck 2018 $\omega_{\text{DE}}=-1.03\pm 0.03$ (Aghanim et al., 2020), and WAMP+CMB $\omega_{\text{DE}}=-1.079^{+0.090}_{-0.089}$ (Hinshaw et al., 2013).
@@
Fig. 5 shows the effective EoS, $\omega_{\text{eff}}$ as a function of redshift $z$ for the $f(\mathcal{G})$ gravity model. At low redshifts ($z\approx 0$), $\omega_{\text{eff}}$ is close to $-1$, indicating DE dominance and accelerated expansion. As $z$ increases, $\omega_{\text{eff}}$ transitions from negative values to less negative values, reflecting a shift from a deceleration phase in the early Universe to an acceleration-dominated phase in the current epoch. The consistent behavior of $\omega_{\text{DE}}$ across both datasets reinforces the capability of the $f(\mathcal{G})$ model to describe the expansion history of the Universe. This transition aligns with theoretical expectations of the $f(\mathcal{G})$ model, which incorporates higher-order curvature terms to account for cosmic dynamics without a cosmological constant. The observed $\omega_{\text{DE}}$ behavior highlights the effectiveness of the model, supporting its viability as an alternative to the $\Lambda$CDM model.
@@
Fig. 6 illustrates the $r-s$ parameter plot for the $f(\mathcal{G})$ gravity model. The trajectory in the $r-s$ plane highlights the evolutionary track of the expansion of the Universe. The $f(\mathcal{G})$ model passes through the region corresponding to the $\Lambda$CDM model, indicated by the red point. At lower values of $s$, the model aligns with quintessence characteristics, suggesting a dynamical DE component with $\omega>-1$. As $s$ increases, the trajectory moves towards regions associated with Chaplygin gas models, indicating a unified dark matter and DE scenario. The smooth transition observed in the $r-s$ parameter space demonstrates the flexibility of the $f(\mathcal{G})$ model in describing different cosmological behaviors. This capability allows the model to account for various dynamics, from quintessence-like to Chaplygin gas-like, providing a comprehensive description of the expansion of the Universe.
@@
Alternatively stated for simplification, when $\text{Om}(z_{1},z_{2})>0$, it indicates quintessence, whereas when $\text{Om}(z_{1},z_{2})<0$, it signifies phantom behavior, where $(z_{1}<z_{2})$. The $\text{Om}(z)$ diagnostic in the $\Lambda$CDM model serves as a null test, as noted in (Sahni et al., 2008), and its sensitivity to the EoS parameter was further explored in subsequent data as seen in (Qi et al., 2018; Zheng et al., 2016; Ding et al., 2015). If $\text{Om}(z)$ remains constant for the redshift, the DE concept would be a cosmological constant. The DE concept will form a cosmological constant if $\text{Om}(z)$ is constant for the redshift. The slope of $\text{Om}(z)$, which is positive for the emerging $\text{Om}(z)$ and denotes phantom phase $(\omega<-1)$ and negative for quintessence region $(\omega>-1)$ also identifies the DE models.
@@
The graph in Fig. 7 shows the reconstructed $\text{Om}(z)$ parameter based on the best-fit data, plotted against redshift. It illustrates a decreasing trend in the $\text{Om}(z)$ parameter as redshift increases. At higher redshifts ($z>1$), the Om diagnostic curve shows a significant decline, indicating the deviation of the $f(\mathcal{G})$ gravity model from the $\Lambda$CDM model. This behavior suggests that the higher-order curvature terms in the $f(\mathcal{G})$ model influence the cosmic dynamics differently compared to the standard cosmological model. The shape of the Om diagnostic curve for the $f(\mathcal{G})$ gravity model highlights its potential to account for cosmic acceleration through modifications to gravity. This diagnostic tool effectively illustrates the differences between the $f(\mathcal{G})$ model and the $\Lambda$CDM model, reinforcing the former’s viability as an alternative explanation for the observed acceleration of the Universe.

### v3

@@ Abstract
In this work, we present a method for numerically solving the Friedmann equations of modified $f(\mathcal{G})$ gravity in the presence of pressureless matter. This method enables us to predict the redshift behaviour of the Hubble expansion rate. To evaluate the credibility of the model, we applied a Bayesian MCMC technique using late-time cosmic observations to impose limitations on the free parameters of the Gauss-Bonnet model. Our results suggest that the $f(\mathcal{G})$ model can reproduce the low-redshift behaviour of the standard Lambda cold dark matter ($\Lambda$CDM) model, but there are significant differences at high redshifts, leading to the absence of a standard matter-dominated epoch. We also examined the profiles of cosmographic parameters using the model parameter values from the standard range to verify the intermediate epochs. Our analysis shows that the highly promising $f(\mathcal{G})$ model is a feasible candidate for explaining the current epochs. We presented a dynamical system analysis framework to examine the stability of the model. Our study identified critical points depicting various phases of the Universe and explained the evolutionary epochs. We demonstrated that the model effectively captures the evolution of energy components over cosmic time, supporting its validity as an alternate explanation for the observed acceleration of the Universe.
@@
cosmology: theory – cosmology: dark energy – cosmological parameters – cosmology: observations.
@@ 1 Introduction
The quest to understand the gravitational force led to the development of general relativity (GR) by Einstein (1915). This theory revolutionized our understanding of gravity by describing it as the curvature of space-time caused by the presence of mass and energy (Misner et al., 1973). Although GR has successfully passed numerous experimental tests (Will, 2014) and remains the cornerstone of modern gravitational physics, it struggles to fully explain phenomena such as the accelerated expansion of the Universe (Riess et al., 1998), dark matter (Bertone et al., 2005), and dark energy (DE) (Copeland et al., 2006). In response to these challenges, several modified theories of gravity have been proposed (Clifton et al., 2012).
@@
In this work, we have explored a subclass of the $f(\mathcal{G})$ model to test its viability as an alternative to the standard cosmological paradigm. We have developed a numerical method to predict the redshift behaviour of the Hubble expansion rate, and our results suggest that the model can reproduce the low-redshift behaviour of the standard Lambda cold dark matter ($\Lambda$CDM) model but has significant differences at high redshifts. The $f(\mathcal{G})$ model is a feasible candidate for explaining the current epochs and effectively captures the evolution of energy components over cosmic time, supporting its validity as an alternative explanation for the observed acceleration of the Universe. We delved into the background cosmological dynamics of the chosen model and evaluated its feasibility using Bayesian analysis supported by Markov Chain Monte Carlo (MCMC) methods applied to late-time cosmic observations, such as Supernovae Ia (Pantheon+) and observational Hubble data (CC sample). We have introduced a dynamical system analysis framework to assess the stability of the model. Our research pinpointed critical points that illustrate different phases of the Universe and elucidated the evolutionary epochs. We have shown that the model effectively represents the changing energy components over cosmic time, supporting its credibility as an alternative explanation for the observed acceleration of the Universe.
This work aims to establish constraints on $f(\mathcal{G})$ cosmology models using CC and Pantheon+ data ses. The paper comprehensively analyses $f(\mathcal{G})$ gravity and uses dynamical system analysis to investigate the stability of the model. The mathematical formalism of $f(\mathcal{G})$ gravity is detailed in Section 2. In Section 3, we use the MCMC method to establish correlations between the $f(\mathcal{G})$ gravity model and observational data to determine the best fits for the model parameters $H_{0}$, $\alpha$, $\beta$, and $m$. Additionally, we present plots of various cosmological parameters such as deceleration, effective equation of state (EoS) $\omega_{\text{eff}}=\frac{p_{\text{eff}}}{\rho_{\text{eff}}}$, Om diagnostic, and the $r-s$ parameter plot, which are essential for understanding the dynamical behaviour of the Universe under $f(\mathcal{G})$ gravity. Subsequently, in Section 4, we construct a dynamical system framework to analyse the critical points of the $f(\mathcal{G})$ gravity model. This analysis is crucial for assessing the stability and viability of the model and its alignment with current cosmological observations. Finally, in Section 5, we present the conclusions of our results.
@@ 2 $R+f(\mathcal{G})$ gravity
The parameter for the effective DE EoS can be defined as follows:
@@ 2.1 Power-Law $f(\mathcal{G})$ Model
In order to determine the theoretical values of the Hubble rate, we can calculate it by solving equation (8a) numerically. When we assume that matter behaves as a pressureless perfect fluid, we can express the matter density as $\rho_{\text{m}}=3H_{0}^{2}\Omega_{\text{m}0}(1+z)^{3}$, with $z$ representing the cosmological redshift (defined as $\frac{a_{0}}{a}=1+z$, where $a_{0}$ represents the scale factor at present and $a$ denotes the scale factor when the light emitted) and $\Omega_{\text{m}0}$ representing the current value of the matter density parameter. Thus, for the particular model we are examining, the first Friedmann equation can be expressed as follows:
@@ 3 Observation with numerical solution
In this part, we will assess the observational feasibility of the model under examination by conducting a Bayesian analysis of the late-time cosmic data. Specifically, we will evaluate the data from the SNe Ia Pantheon+ sample (Brout et al., 2022) and the cosmic chronometers (CC) derived from the observational Hubble data compiled in Ref. (Moresco et al., 2022). We have not assumed that the Hubble and Pantheon data sets are correlated. Rather, we will present our results independently for the CC and Pantheon+ data sets. Utilizing these data sets for statistical analysis enables us to obtain reliable results that are not influenced by assumptions of any particular underlying reference model (D’Agostino & Luongo, 2018; D’Agostino, 2019; Lohakare et al., 2023). In the following subsections, we will outline the key characteristics of these measurements and the corresponding likelihood functions.
@@ 3.1 Cosmic Chronometers (CC)
Figure 1: Contour plots show the $1\sigma$ and $2\sigma$ uncertainty regions for the variables $H_{0}$, $\Omega_{m0}$, $\alpha$, $\beta$, and $m$. These contours are derived from the CC sample (left panel) and the Pantheon+ data (right panel).
@@ 3.2 Supernovae type Ia (SNe Ia)
The contour plot for the parameters $H_{0}$ and $\Omega_{\text{m}0}$ exhibits a more elliptical shape compared to the other parameter pairs, which display more square or elongated contours. This elliptical shape indicates that there is a relatively weak correlation between $H_{0}$ and $\Omega_{\text{m}0}$, suggesting that the data independently constrains these two parameters. This independence implies that the variations in one parameter do not significantly affect the value of the other, leading to a more symmetrical uncertainty region. In contrast, the square or elongated contours seen in other parameter pairs indicate stronger correlations, where changes in one parameter can be offset by adjustments in another to maintain a similar fit to the data. This strong coupling results in less symmetrical and more stretched uncertainty regions, reflecting the interdependence of these parameters within the $f(\mathcal{G})$ gravity model. The best-fit values derived from the MCMC analysis are presented in Table 1. In order to evaluate the efficacy of our MCMC analysis, we calculated the associated AIC, $\text{AIC}_{\text{c}}$ and BIC values, which are presented in Table 2. Our findings strongly support the assumed $f(\mathcal{G})$ gravity models when analysing the data sets. Moreover, we noted that the $f(\mathcal{G})$ model demonstrates greater precision when applied to the Pantheon+ data sets.
@@
Figure 3: Whisker plot depicting the model parameters $H_{0}$, $\Omega_{\text{m}0}$, $\alpha$, $\beta$ and $m$, respectively, highlights their discrepancies.
@@ 3.3 Cosmological Parameter Evolution
We analyse the evolution of crucial cosmological parameters, including the effective EoS, statefinder, and Om diagnostic parameters, by imposing constraints on model parameters using various observational data. We present a fully general expression for the deceleration parameter, $q=-\ddot{a}/aH^{2}$, as follows:
@@
Figure 4: Graphical representation of the deceleration parameter versus redshift using the constrained coefficients from Fig. 1. The thick line represents the behaviour of the deceleration parameter for the $f(\mathcal{G})$ model, while the dashed line shows the deceleration parameter for the $\Lambda$CDM model.
@@
Fig. 4 presents the deceleration parameter as a function of redshift $z$ for the $f(\mathcal{G})$ gravity model, derived from both Hubble data (upper panel) and Pantheon+ data (lower panel). Fig. 4 demonstrate that the restricted values of model parameters derived from the analysed CC and Pantheon+ data sets indicate a transition of $q$ from positive (indicating early deceleration) in the past to negative (indicating current acceleration) in the present. The present value of deceleration parameter $q_{0}$ is measured to be $-0.527$ and $-0.499$ for the CC and Pantheon+ data, respectively, at the current cosmic epoch, which aligns relatively well with the range of $q_{0}=-0.528^{+0.092}_{-0.088}$ determined by recent observations (Gruber & Luongo, 2014). Recent observations are consistent with this deceleration parameter, and the resulting model indicates a smooth transition from deceleration to acceleration at $z_{t}=0.84$ and $z_{t}=0.82$ for the CC and Pantheon+ data sets, respectively. The derived transition redshift $z_{t}$ aligns with current constraints based on 11 $H(z)$ observations reported by Busca et al. (2013) for redshifts $0.2\leq z\leq 2.3$, $z_{t}=0.74\pm 0.5$ from Farooq & Ratra (2013), $z_{t}=0.7679^{+0.1831}_{-0.1829}$ by Capozziello et al. (2014), and $z_{t}=0.60^{+0.21}_{-0.12}$ by Yang & Gong (2020). The consistency between the curves from Hubble and Pantheon+ data underscores the robustness of the $f(\mathcal{G})$ model in capturing the expansion dynamics of the Universe.
@@
Figure 5: Graphical representation of the EoS parameter versus redshift using the constrained coefficients from Fig. 1. The thick line represents the behaviour of the EoS parameter for the $f(\mathcal{G})$ model, while the dashed line shows the EoS parameter for the $\Lambda$CDM model.
@@
The deceleration parameter is one of the key factors that characterize the behaviour of the Universe, determining whether it continuously decelerates, accelerates, or undergoes multiple phases of transition. Similarly, energy sources influence the evolution of the Universe through the EoS parameter, defined as $\omega_{\text{DE}}$ is shown in Fig. 5. By calculating the energy density and pressure of DE, as depicted in Fig. 5, we can observe the variations in the effective EoS of DE relative to the redshift variable. The current EoS values for DE, $\omega_{\text{DE}}(z=0)$, are obtained as $-1.018$, $-0.999$ for the CC, Pantheon+, data sets, respectively. These values indicate phantom behaviour (at $z<0$) and a trend towards approximately $-1.32$ at late times. The present values of $\omega_{\text{eff}}$ are $-0.684$, $-0.666$ for the CC and Pantheon+ data sets, respectively. Various cosmological studies have also constrained the EoS parameter, including the Supernovae Cosmology Project $\omega_{\text{DE}}=-1.035^{+0.055}_{-0.059}$ (Amanullah et al., 2010), Planck 2018 $\omega_{\text{DE}}=-1.03\pm 0.03$ (Aghanim et al., 2020), and WAMP+CMB $\omega_{\text{DE}}=-1.079^{+0.090}_{-0.089}$ (Hinshaw et al., 2013).
@@
Fig. 5 shows the effective EoS, $\omega_{\text{eff}}$ as a function of redshift $z$ for the $f(\mathcal{G})$ gravity model. At low redshifts ($z\approx 0$), $\omega_{\text{eff}}$ is close to $-1$, indicating DE dominance and accelerated expansion. As $z$ increases, $\omega_{\text{eff}}$ transitions from negative values to less negative values, reflecting a shift from a deceleration phase in the early Universe to an acceleration-dominated phase in the current epoch. The consistent behaviour of $\omega_{\text{DE}}$ across both data sets reinforces the capability of the $f(\mathcal{G})$ model to describe the expansion history of the Universe. This transition aligns with theoretical expectations of the $f(\mathcal{G})$ model, which incorporates higher-order curvature terms to account for cosmic dynamics without a cosmological constant. The observed $\omega_{\text{DE}}$ behaviour highlights the effectiveness of the model, supporting its viability as an alternative to the $\Lambda$CDM model.
@@
Fig. 6 illustrates the $r-s$ parameter plot for the $f(\mathcal{G})$ gravity model. The trajectory in the $r-s$ plane highlights the evolutionary track of the expansion of the Universe. The $f(\mathcal{G})$ model passes through the region corresponding to the $\Lambda$CDM model, indicated by the red point. At lower values of $s$, the model aligns with quintessence characteristics, suggesting a dynamical DE component with $\omega>-1$. As $s$ increases, the trajectory moves towards regions associated with Chaplygin gas models, indicating a unified dark matter and DE scenario. The smooth transition observed in the $r-s$ parameter space demonstrates the flexibility of the $f(\mathcal{G})$ model in describing different cosmological behaviours. This capability allows the model to account for various dynamics, from quintessence-like to Chaplygin gas-like, providing a comprehensive description of the expansion of the Universe.
@@
Alternatively stated for simplification, when $\text{Om}(z_{1},z_{2})>0$, it indicates quintessence, whereas when $\text{Om}(z_{1},z_{2})<0$, it signifies phantom behaviour, where $(z_{1}<z_{2})$. The $\text{Om}(z)$ diagnostic in the $\Lambda$CDM model serves as a null test, as noted in (Sahni et al., 2008), and its sensitivity to the EoS parameter was further explored in subsequent data as seen in (Qi et al., 2018; Zheng et al., 2016; Ding et al., 2015). If $\text{Om}(z)$ remains constant for the redshift, the DE concept would be a cosmological constant. The DE concept will form a cosmological constant if $\text{Om}(z)$ is constant for the redshift. The slope of $\text{Om}(z)$, which is positive for the emerging $\text{Om}(z)$ and denotes phantom phase $(\omega<-1)$ and negative for quintessence region $(\omega>-1)$ also identifies the DE models.
@@
The graph in Fig. 7 shows the reconstructed $\text{Om}(z)$ parameter based on the best-fitting data, plotted against redshift. It illustrates a decreasing trend in the $\text{Om}(z)$ parameter as redshift increases. At higher redshifts ($z>1$), the Om diagnostic curve shows a significant decline, indicating the deviation of the $f(\mathcal{G})$ gravity model from the $\Lambda$CDM model. This behaviour suggests that the higher-order curvature terms in the $f(\mathcal{G})$ model influence the cosmic dynamics differently compared to the standard cosmological model. The shape of the Om diagnostic curve for the $f(\mathcal{G})$ gravity model highlights its potential to account for cosmic acceleration through modifications to gravity. This diagnostic tool effectively illustrates the differences between the $f(\mathcal{G})$ model and the $\Lambda$CDM model, reinforcing the former’s viability as an alternative explanation for the observed acceleration of the Universe.